two positive numbers are in the ratio of 5:12. if there product is 960 what is the positive diffrence between the two numbers
Answer by edjones(8007)

Let x=a number that when multiplied by both 5 and 12 makes the resulting product 960.
5x*12x=960
60x^2=960
x^2=16
x=4
5*4=20
12*4=48
48*20=960
.
48 and 20 are the two numbers.
.
Ed